### v1

@@ 4.4.2 Key propositions
$K_{\theta}:=\Big\lfloor\frac{\lfloor c_{0}n\rfloor}{\lfloor\theta n\rfloor}\Big\rfloor,\qquad h_{r}:=\bar{h}+\lfloor c_{0}n\rfloor-r\lfloor\theta n\rfloor,\quad 0\leq r\leq K_{\theta},$ (68)
@@
Clearly,
$\mathcal{N}_{i}\;\geq\;\sum_{r=1}^{K_{\theta}}\mathcal{N}_{i,r},$
with equality when $\lfloor c_{0}n\rfloor/\lfloor\theta n\rfloor\in{\mathbb{N}}$.
@@
First, for any vertex $v$ and $0\leq j\leq|v|$, denote by $\text{Anc}(v,j)$ the ancestor of $v$ at height $j$. For $0\leq r\leq K_{\theta}$ and $1\leq i\leq m$, denote the ancestral lineage of $u^{h_{r-1}}_{i}$ up to height $h_{r}$ by
$U_{i,r}:=U_{i}\big([h_{r-1},h_{r}]\big):=\Big(u^{h_{r-1}}_{i},\text{Anc}(u^{h_{r-1}}_{i},h_{r-1}-1),\ldots,\text{Anc}(u^{h_{r-1}}_{i},h_{r})\Big).$
For each $j$ with $h_{r}<j\leq h_{r-1}$, denote the remaining $q_{n}(j)-m$ vertices at height $j$ by $v^{j}_{1},\ldots,v^{j}_{q_{n}(j)-m}$, where the subscripts are assigned, conditionally on $\mathscr{T}^{n}$, uniformly at random and independently across heights. (Unlike $(y^{j}_{l})$ before, we specify the distribution of subscripts to simplify later analysis.)
Then, for $1\leq i\leq m$ and $1\leq r\leq K_{\theta}$, set
@@
For $0\leq r\leq K_{\theta}$, we define the event
@@
$\mathbb{Q}^{m}_{n}\big(f(\mathcal{N}_{i,r})\big)=\mathbb{P}_{n}\left(f(\widetilde{\mathcal{N}}_{i,r})\,|\,\widetilde{B}_{r}\right)$ (69)
@@
Assume (H1) and (H2’). Then there exists constant $C_{0}\in(0,1)$ such that for any $\theta\in(0,c_{0})$ sufficiently small, there exists $\gamma=\gamma(\theta,C_{0})>0$ satisfying
@@
The proofs of Propositions 4.11 and 4.12 are postponed, where we will see what additional condition $\theta$ should satisfy. We first explain how these results imply Theorem 4.9.
Proof 4.13 (Proof of Theorem 4.9 assuming Propositions 4.11 and 4.12).
Observe that if $\mathcal{N}_{i,r}\geq\gamma n^{2}$ for some $1\leq r\leq K_{\theta}$, then $\mathcal{N}_{i}\geq\gamma n^{2}$. Hence, by Propositions 4.11 and 4.12, for any $\theta>0$, there exists $\gamma>0$ such that for sufficiently large $n$, $1\leq i\leq m\leq M_{0}$ and $0\leq r\leq K_{\theta}-1$,
${\mathbb{Q}}_{n}^{m}(\mathcal{N}_{i}\geq\gamma n^{2})\;\geq\;1-{\mathbb{Q}}_{n}^{m}\!\big(\mathcal{N}_{i,r}<\gamma n^{2},\ \forall\,1\leq r\leq K_{\theta}\big)\;\geq\;1-(1-C_{0})^{K_{\theta}}.$
Since $K_{\theta}\sim c_{0}/\theta$ as $n\to\infty$, for any $\eta>0$ one can choose $\theta=\theta(\eta)>0$ sufficiently small such that, for all $n\geq 1$,
$(1-C_{0})^{K_{\theta}}\leq\frac{\eta}{2M_{0}}.$
Thus, there exists $\gamma>0$ such that
[…]
Consequently,
@@
which completes the proof.
4.4.3 Proof of Propositions 4.11 and 4.12
We begin with a preliminary lemma, which will be used repeatedly in the sequel. Recall the definitions of $\mathcal{G}_{j}$ in (27) and labeling map $\Lambda_{j}$ in (66).
@@
Fix a base height $h^{*}$. For $0\leq h\leq h^{*}$ and $1\leq k\leq q_{n}(h)$, let $(A,\mathcal{V})$ be a pair consisting of
  • an event $A$ measurable with respect to $\mathcal{G}_{h}\vee\sigma(\Lambda_{h})$, and
@@
We say that $(A,\mathcal{V})$ is a $(h,k)$-uniform selection pair if, under $\mathbb{P}_{n}$,
@@
that is, conditionally on $A$, the vector $\mathcal{V}$ has the same distribution as a uniformly chosen $k$-tuple of distinct vertices at height $h$.
@@
Let $0\leq h\leq h^{*}$ and $1\leq k\leq q_{n}(h)$, and let $(A,\mathcal{V})$ be a $(h,k)$-uniform selection pair. Then for any test function $f$ defined on the value space of $\big((u^{h}_{i})_{i=1,\ldots,k},(\nu^{i})_{1\leq i\leq h-1}\big)$, we have
$\mathbb{P}_{n}\Big[f\big(\mathcal{V},(\nu^{i})_{1\leq i\leq h-1}\big)\,\big|\,A\Big]=\mathbb{P}_{n}\Big[f\big((u^{h}_{i})_{1\leq i\leq k},(\nu^{i})_{1\leq i\leq h-1}\big)\Big].$
@@
Proof 4.16 (Proof of Proposition 4.11).
Recall the definition of $E(n,\delta,\eta,m)$ in (65). For each $0\leq r\leq K_{\theta}$, we split the event $E(n,\delta,\eta,m)$ according to the behavior of the process above and below $h_{r}$:
$E(n,\delta,\eta,m)=A_{r}\cap B_{r},$
@@
Then for any $1\leq r\leq K_{\theta}$ and subsets $D_{1},\ldots,D_{r}$ of $\mathbb{N}$, we have
$\mathbb{Q}^{m}_{n}\Bigg[\bigcap_{l=1}^{r}\{\mathcal{N}_{i,l}\in D_{l}\}\Bigg]=\frac{\mathbb{P}_{n}\Big[A_{r}\cap B_{r}\cap\bigcap_{l=1}^{r}\{\mathcal{N}_{i,l}\in D_{l}\}\Big]}{\mathbb{P}_{n}(A_{r}\cap B_{r})}.$ (70)
@@
Proof 4.18.
Clearly, both $A_{r}$ and $\bigcap_{l=1}^{r}\{\mathcal{N}_{i,l}\in D_{l}\}$ are measurable with respect to $\mathcal{G}_{h_{r}}\vee\sigma(\Lambda_{h_{r}})$. Furthermore, as in the proof of Proposition 3.12, the uniformity condition follows directly from the exchangeability of $\big(\mathscr{T}^{n}(v):|v|=h_{r}\big)$. This concludes the lemma.
Returning to (70), the next step is to apply Lemma 4.15 to the numerator on the RHS. Recall the definitions of $\widetilde{B}_{r}$ and $\widetilde{\mathcal{N}}_{i,r}$ introduced before Proposition 4.11.
By Lemmas 4.15 and 4.17, for $1\leq r\leq K_{\theta}$,
$\mathbb{P}_{n}\Big[A_{r}\cap B_{r}\cap\bigcap_{l=1}^{r}\{\mathcal{N}_{i,l}\in D_{l}\}\Big]$ $=\mathbb{P}_{n}\Big[A_{r-1}\cap\bigcap_{l=1}^{r-1}\{\mathcal{N}_{i,l}\in D_{l}\}\Big]\,\mathbb{P}_{n}\big(\widetilde{B}_{r}\cap\{\widetilde{\mathcal{N}}_{i,r}\in D_{r}\}\big),$
$\mathbb{P}_{n}\Big[A_{r}\cap B_{r}\cap\bigcap_{l=1}^{r-1}\{\mathcal{N}_{i,l}\in D_{l}\}\Big]$ $=\mathbb{P}_{n}\Big[A_{r-1}\cap\bigcap_{l=1}^{r-1}\{\mathcal{N}_{i,l}\in D_{l}\}\Big]\,\mathbb{P}_{n}(\widetilde{B}_{r}).$
@@
$\mathbb{P}_{n}\Big[A_{r}\cap B_{r}\cap\bigcap_{l=1}^{r}\{\mathcal{N}_{i,l}\in D_{l}\}\Big]=\mathbb{P}_{n}\Big[A_{r}\cap B_{r}\cap\bigcap_{l=1}^{r-1}\{\mathcal{N}_{i,l}\in D_{l}\}\Big]\,\mathbb{P}_{n}\big(\widetilde{\mathcal{N}}_{i,r}\in D_{r}\,\big|\,\widetilde{B}_{r}\big).$
Substituting into (70), we obtain
@@
which proves the desired results.
Next, we turn to the proof of Proposition 4.12. By (69), it suffices to show that
@@
with the convention $\inf\emptyset=\infty$. In words, $\tau_{i}(v^{j}_{l})$ is the “first time” $k$ at which the ancestral lineage of $v^{j}_{l}$ merges into the trunk path $U_{i,r}$. Clearly,
@@
    ${\mathbb{P}}_{n}\big(\tau_{i}(v^{j}_{l})<\infty\,\big|\,\widetilde{B}_{r}\big)=\frac{1}{m}\left\{1-\prod_{h=0}^{j-h_{r}-1}\Bigg[1-\frac{m\sigma^{2}_{n}(j-h)}{q_{n}(j-h)}\big(1+o_{n}(1)\big)\Bigg]\right\}.$ (71)
  • (2)
@@
    ${\mathbb{P}}_{n}\big(\tau_{i}(v^{j}_{l})<\infty\,\big|\,\tau_{i}(v^{j^{\prime}}_{l^{\prime}})<\infty,\,\widetilde{B}_{r}\big)$ (72)
@@
Proof 4.20 (Proof of Proposition 4.12 assuming Proposition 4.19).
@@
$\mathbb{E}_{n}\big(\widetilde{\mathcal{N}}_{i,r}\,\big|\,\widetilde{B}_{r}\big)$ $=\sum_{j=h_{r}+1}^{h_{r-1}}\sum_{l=1}^{q_{n}(j)-m}{\mathbb{P}}_{n}\big(\tau_{i}(v^{j}_{l})<\infty\,\big|\,\widetilde{B}_{r}\big).$ (73)
We notice that $j-h_{r}\leq\theta n$ for $h_{r}<j\leq h_{r-1}$. By (71), (H1) and (H2’), we deduce that for all $n$ sufficiently large,
@@
provided that $\theta\leq\frac{1}{2c_{1}M_{0}}\log 2$ (in which case $\frac{c_{1}m}{n}\leq\frac{\log 2}{2\theta n}\leq 1-(\frac{1}{2})^{\frac{1}{\theta n}}$ for all $n$ sufficiently large); the last equality then follows since $(1-x)^{k}\leq 1-\frac{k}{2}x$ for all $0\leq x\leq 1-(\frac{1}{2})^{\frac{1}{\theta n}}$ and $1\leq k\leq\theta n$. Substituting this into (73) gives
$\mathbb{E}_{n}\big(\widetilde{\mathcal{N}}_{i,r}\,\big|\,\widetilde{B}_{r}\big)\geq\sum_{j=h_{r}+1}^{h_{r-1}}c_{2}n\cdot\frac{c_{1}(j-h_{r})}{2n}=\frac{c_{1}c_{2}}{4}\lfloor\theta n\rfloor(\lfloor\theta n\rfloor-1)\geq c_{3}\theta^{2}n^{2}.$ (74)
@@
$\begin{aligned} \mathbb{E}_{n}\big((\widetilde{\mathcal{N}}_{i,r})^{2}\,\big|\,\widetilde{B}_{r}\big)&=\mathbb{E}_{n}\big(\widetilde{\mathcal{N}}_{i,r}\,\big|\,\widetilde{B}_{r}\big)+\sum_{(j,l)\neq(j^{\prime},l^{\prime})}\mathbb{P}_{n}\big(\tau_{i}(y^{j}_{l}),\,\tau_{i}(y^{j^{\prime}}_{l^{\prime}})<\infty\,\big|\,\widetilde{B}_{r}\big)\\ &\leq\mathbb{E}_{n}\big(\widetilde{\mathcal{N}}_{i,r}\,\big|\,\widetilde{B}_{r}\big)+2\sum_{(j,l)\neq(j^{\prime},l^{\prime}):j\leq j^{\prime}}\mathbb{P}_{n}\big(\tau_{i}(y^{j}_{l}),\,\tau_{i}(y^{j^{\prime}}_{l^{\prime}})<\infty\,\big|\,\widetilde{B}_{r}\big),\end{aligned}$ (75)
@@
By (79), (H1) and (H2’), for $(j,l)\neq(j^{\prime},l^{\prime})$ with $j\leq j^{\prime}$,
${\mathbb{P}}_{n}\big(\tau_{i}(v^{j}_{l})<\infty\,\big|\,\tau_{i}(v^{j^{\prime}}_{l^{\prime}})<\infty,\,\widetilde{B}_{r}\big)$ $\leq\frac{1}{m}\left\{1-\big(1+o_{n}(1)\big)\Bigg[1-\frac{c_{4}m}{n}\Bigg]^{\theta n}\right\}\leq c_{5}\theta,$
@@
$\leq c_{5}\theta\cdot\sum_{(j,l)\neq(j^{\prime},l^{\prime}):j\leq j^{\prime}}\mathbb{P}_{n}\big(\tau_{i}(y^{j^{\prime}}_{l^{\prime}})<\infty\,\big|\,\widetilde{B}_{r}\big)$
$\leq c_{6}\theta^{2}n^{2}\cdot\mathbb{E}_{n}\big(\widetilde{\mathcal{N}}_{i,r}\,\big|\,\widetilde{B}_{r}\big).$
Substituting this estimate into (75) yields
@@
$\,\ \leq c_{7}\Big[\mathbb{E}_{n}\big(\widetilde{\mathcal{N}}_{i,r}\,\big|\,\widetilde{B}_{r}\big)\Big]^{2},$
for all large $n$. Finally, taking $\gamma=\tfrac{1}{2}c_{3}\theta^{2}$, we obtain
@@
This completes the proof.
@@
$\mathbb{P}_{n}\big(\tau(v^{j}_{l})=\infty\,\big|\,\widetilde{B}_{r}\big)=\prod_{h=0}^{j-h_{r}-1}\Bigg[1-\frac{m\sigma^{2}_{n}(j-h)}{q_{n}(j-h)}\big(1+o_{n}(1)\big)\Bigg].$ (76)
Assuming Lemma 4.22, the desired estimate (71) follows directly from the exchangeability of the vector $\big(\mathscr{T}^{n}(v):|v|=h_{r}\big)$.
Proof 4.23 (Proof of Lemma 4.22).
@@
$\mathbb{P}_{n}\big(\tau(v^{j}_{l})>h+1\,\big|\,\tau(v^{j}_{l})>h,\,\widetilde{B}_{r}\big)=1-\frac{m\sigma^{2}_{n}(j-h)}{q_{n}(j-h)}\big(1+o_{n}(1)\big).$ (77)
@@
Then we have
$\widetilde{B}_{r}=\widetilde{B}^{[j-h,h_{r-1}]}_{r}\cap\widetilde{B}^{[h_{r},j-h]}_{r}$.
By Proposition 3.12 and the uniform ordering of vertices $v^{j}_{l}$, the vector
@@
together with the event $\{\tau(v^{j}_{l})>h\}\cap\widetilde{B}^{[j-h,h_{r-1}]}_{r}$, forms a (jh,m+1)(j-h,m+1)-uniform selection pair, with $h_{r-1}$ serving as the base height $h^{*}$ in Definition 4.14. Hence, by Lemma 4.17,
$\begin{aligned} \mathbb{P}_{n}\big(\tau(v^{j}_{l})>h+1,\,\widetilde{B}_{r}\big)&=\mathbb{P}_{n}\big(\tau(v^{j}_{l})>h,\,\widetilde{B}^{[j-h,h_{r-1}]}_{r}\big)\\ &\quad\cdot\mathbb{P}_{n}\Big[\mathfrak{p}(u^{j-h}_{1}),\ldots,\mathfrak{p}(u^{j-h}_{m+1})\text{ are distinct,}\\ &\quad\quad\quad\quad\#\big\{\text{Anc}(u^{j-h}_{i},h_{r}):1\leq i\leq m\big\}=m\Big].\end{aligned}$ (78)
@@
$\Big(\big\{\mathfrak{p}(u^{j-h}_{j}):1\leq i\leq m\big\},\ \{\mathfrak{p}(u^{j-h}_{1}),\ldots,\mathfrak{p}(u^{j-h}_{m+1})\text{ are distinct}\}\Big)$
is a $(1,m)$-uniform selection pair with jhj-h as the base height. Applying Lemma 4.17 again yields

### v2

@@ 4.4.2 Key propositions
$K_{\theta}:=\bigg\lfloor\frac{\lfloor c_{0}n\rfloor}{\lfloor\theta n\rfloor}\bigg\rfloor,\qquad h_{r}:=\bar{h}+\lfloor c_{0}n\rfloor-r\lfloor\theta n\rfloor,\quad 0\leq r\leq K_{\theta},$ (4.11)
@@
Clearly, $\mathcal{N}_{i}\;\geq\;\sum_{r=1}^{K_{\theta}}\mathcal{N}_{i,r}\,$, with equality when $\lfloor c_{0}n\rfloor/\lfloor\theta n\rfloor\in{\mathbb{N}}$.
@@
Recall the notion of $\text{Anc}(v,j)$ appearing at the beginning of Section 3.2, which represents the ancestor of $v$ at height $j$. For $1\leq r\leq K_{\theta}$ and $1\leq i\leq m$, denote the ancestral lineage of $u^{h_{r-1}}_{i}$ up to height $h_{r}$ by
$U_{i,r}:=\Big(u^{h_{r-1}}_{i},\text{Anc}(u^{h_{r-1}}_{i},h_{r-1}-1),\ldots,\text{Anc}(u^{h_{r-1}}_{i},h_{r})\Big).$
For each $j$ with $h_{r}<j\leq h_{r-1}$, denote the remaining $q_{n}(j)-m$ vertices at height $j$ by $v^{j}_{1},\ldots,v^{j}_{q_{n}(j)-m}$, where the subscripts are assigned, conditionally on $\mathscr{T}^{n}$, uniformly at random and independently across heights.  [^5] Unlike $(y^{j}_{l})$ before, we specify the distribution of subscripts to simplify later analysis. Then we set, for $1\leq r\leq K_{\theta}$ and $1\leq i\leq m$,
@@
For $1\leq r\leq K_{\theta}$, we define the event
@@
$\mathbb{Q}^{m}_{n}\big(f(\mathcal{N}_{i,r})\big)=\mathbb{P}_{n}\left(f(\widetilde{\mathcal{N}}_{i,r})\,|\,\widetilde{B}_{r}\right)$ (4.12)
@@
Assume (H1)–(H2). Then there exists constant $C_{0}\in(0,1)$ such that for any $\theta\in(0,c_{0})$ sufficiently small, there exists $\gamma=\gamma(\theta,C_{0})>0$ satisfying
@@
The proofs of Propositions 4.9 and 4.10 are postponed. We first explain how these results imply Theorem 4.8.
Proof of Theorem 4.8 assuming Propositions 4.9 and 4.10.
Observe that if $\mathcal{N}_{i,r}\geq\gamma n^{2}$ for some $1\leq r\leq K_{\theta}$, then $\mathcal{N}_{i}\geq\gamma n^{2}$. Hence, by Propositions 4.9 and 4.10, for any sufficiently small $\theta\in(0,c_{0})$, there exists $\gamma>0$ such that for sufficiently large $n$, $1\leq i\leq m\leq M_{0}$ and $0\leq r\leq K_{\theta}-1$,
${\mathbb{Q}}_{n}^{m}(\mathcal{N}_{i}\geq\gamma n^{2})\;\geq\;1-{\mathbb{Q}}_{n}^{m}\!\big(\mathcal{N}_{i,r}<\gamma n^{2},\ \forall\,1\leq r\leq K_{\theta}\big)\;\geq\;1-(1-C_{0})^{K_{\theta}}.$ (4.13)
On the other hand, since $K_{\theta}\sim c_{0}/\theta$ as $n\to\infty$, for any $\eta>0$ one can choose $\theta=\theta(\eta)>0$ small enough such that
$(1-C_{0})^{K_{\theta}}\leq\frac{\eta}{2M_{0}},\qquad\text{for all }n\geq 1.$ (4.14)
Combining (4.13) and (4.14), we obtain that for any $\eta>0$, there exists $\gamma>0$ such that for all sufficiently large $n$,
@@
which completes the proof. ∎
4.4.3 Proof of Propositions 4.9 and 4.10
We begin with a preliminary lemma, which will be used repeatedly in the sequel. Recall the definitions of $\mathcal{G}^{(k)}_{j}$ in (3.18) and labeling map $\Lambda_{j}$ in (4.9).
@@
Fix a base height $h^{*}$. For $0\leq h\leq h^{*}$, $1\leq p\leq q_{n}(h^{*})$ and $1\leq k\leq q_{n}(h)$, let $(A,\mathcal{V})$ be a pair consisting of
  • an event $A$ measurable with respect to $\mathcal{G}_{h}^{(p)}\vee\sigma(\Lambda_{h})$, and
@@
We say that $(A,\mathcal{V})$ is a $(p,h,k)$-uniform selection pair if, under $\mathbb{P}_{n}$,
@@
that is, conditionally on $A$, the vector $\mathcal{V}$ has the same distribution as a uniformly chosen $k$-tuple of distinct vertices at height $h$. When $p=q_{n}(h^{*})$, we abbreviate this as an $(h,k)$-uniform selection pair.
@@
Let $0\leq h\leq h^{*}$, $1\leq p\leq q_{n}(h^{*})$ and $1\leq k\leq q_{n}(h)$, and let $(A,\mathcal{V})$ be a $(p,h,k)$-uniform selection pair. Then for any test function $f$ defined on the value space of $\big((u^{h}_{i})_{i=1,\ldots,k},(\nu^{s})_{1\leq s\leq h-1}\big)$, we have
$\mathbb{P}_{n}\Big[f\big(\mathcal{V},(\nu^{s})_{1\leq s\leq h-1}\big)\,\big|\,A\Big]=\mathbb{P}_{n}\Big[f\big((u^{h}_{i})_{1\leq i\leq k},(\nu^{s})_{1\leq s\leq h-1}\big)\Big].$
@@
Proof of Proposition 4.9.
Recall the definition of $E(n,\delta,m)$ in (4.8). For each $0\leq r\leq K_{\theta}$, we split the event $E(n,\delta,m)$ according to the behavior of the process above and below $h_{r}$:
$E(n,\delta,m)=A_{r}\cap B_{r},$
@@
Then for any $1\leq r,r^{\prime}\leq K_{\theta}$ and subsets $a_{1},\ldots,D_{r}$ of $\mathbb{N}$, we have
$\mathbb{Q}^{m}_{n}\Bigg[\bigcap_{l=1}^{r^{\prime}}\{\mathcal{N}_{i,l}\in D_{l}\}\Bigg]=\frac{\mathbb{P}_{n}\Big[A_{r}\cap B_{r}\cap\bigcap_{l=1}^{r^{\prime}}\{\mathcal{N}_{i,l}\in D_{l}\}\Big]}{\mathbb{P}_{n}(E(n,\delta,m))}.$ (4.15)
@@
Proof.
Clearly, both $A_{r}$ and $\bigcap_{l=1}^{r}\{\mathcal{N}_{i,l}\in D_{l}\}$ are measurable with respect to $\mathcal{G}_{h_{r}}\vee\sigma(\Lambda_{h_{r}})$. Furthermore, as in the proof of Proposition 3.8, the uniformity condition follows directly from the exchangeability of $\big(\mathscr{T}^{n}(v):|v|=h_{r}\big)$. This concludes the lemma. ∎
Returning to (4.15), the next step is to apply Lemma 4.12 to the numerator on the RHS. Recall the definitions of $\widetilde{B}_{r}$ and $\widetilde{\mathcal{N}}_{i,r}$ introduced before Proposition 4.9. By Lemmas 4.12 and 4.13, for $1\leq r\leq K_{\theta}$,
$\mathbb{P}_{n}\Big[A_{r}\cap B_{r}\cap\bigcap_{l=1}^{r+1}\{\mathcal{N}_{i,l}\in D_{l}\}\Big]$ $=\mathbb{P}_{n}\Big[A_{r}\cap\bigcap_{l=1}^{r}\{\mathcal{N}_{i,l}\in D_{l}\}\Big]\,\mathbb{P}_{n}\big(\widetilde{B}_{r}\cap\{\widetilde{\mathcal{N}}_{i,r+1}\in D_{r}\}\big),$
$\mathbb{P}_{n}\Big[A_{r}\cap B_{r}\cap\bigcap_{l=1}^{r}\{\mathcal{N}_{i,l}\in D_{l}\}\Big]$ $=\mathbb{P}_{n}\Big[A_{r}\cap\bigcap_{l=1}^{r}\{\mathcal{N}_{i,l}\in D_{l}\}\Big]\,\mathbb{P}_{n}(\widetilde{B}_{r}).$
@@
$\mathbb{P}_{n}\Big[A_{r}\cap B_{r}\cap\bigcap_{l=1}^{r+1}\{\mathcal{N}_{i,l}\in D_{l}\}\Big]=\mathbb{P}_{n}\Big[A_{r}\cap B_{r}\cap\bigcap_{l=1}^{r}\{\mathcal{N}_{i,l}\in D_{l}\}\Big]\,\mathbb{P}_{n}\big(\widetilde{\mathcal{N}}_{i,r+1}\in D_{r}\,\big|\,\widetilde{B}_{r}\big).$
Substituting into (4.15), we obtain
@@
which proves the desired results. ∎
Next, we turn to the proof of Proposition 4.10. By (4.12), it suffices to show that
@@
with the convention $\inf\emptyset=\infty$. In other words, $\tau_{i}(v^{j}_{l})$ is the “first time” $k$ at which the ancestral lineage of $v^{j}_{l}$ merges into the trunk path $U_{i,r}$. Clearly,
@@
    ${\mathbb{P}}_{n}\big(\tau_{i}(v^{j}_{l})<\infty\,\big|\,\widetilde{B}_{r}\big)=\frac{1}{m}\left\{1-\prod_{h=0}^{j-h_{r}-1}\Bigg[1-\frac{m\sigma^{2}_{n}(j-h)}{q_{n}(j-h)}\big(1+o_{n}(1)\big)\Bigg]\right\}.$ (4.16)
  2. (2)
@@
    ${\mathbb{P}}_{n}\big(\tau_{i}(v^{j}_{l})<\infty\,\big|\,\tau_{i}(v^{j^{\prime}}_{l^{\prime}})<\infty,\,\widetilde{B}_{r}\big)$ (4.17)
@@
Proof of Proposition 4.10 assuming Proposition 4.14.
@@
$\mathbb{E}_{n}\big(\widetilde{\mathcal{N}}_{i,r}\,\big|\,\widetilde{B}_{r}\big)$ $=\sum_{j=h_{r}+1}^{h_{r-1}}\sum_{l=1}^{q_{n}(j)-m}{\mathbb{P}}_{n}\big(\tau_{i}(v^{j}_{l})<\infty\,\big|\,\widetilde{B}_{r}\big).$ (4.18)
We notice that $j-h_{r}\leq\theta n$ for $h_{r}<j\leq h_{r-1}$. By (4.16), (H1) and (H2), we deduce that for all $n$ sufficiently large,
@@
provided that $\theta\leq\frac{1}{2c_{1}M_{0}}\log 2$ (in which case $\frac{c_{1}m}{n}\leq\frac{\log 2}{2\theta n}\leq 1-(\frac{1}{2})^{\frac{1}{\theta n}}$ for all $n$ sufficiently large); the last equality then follows since $(1-x)^{k}\leq 1-\frac{k}{2}x$ for all $0\leq x\leq 1-(\frac{1}{2})^{\frac{1}{\theta n}}$ and $1\leq k\leq\theta n$. Substituting this into (4.18) gives
$\mathbb{E}_{n}\big(\widetilde{\mathcal{N}}_{i,r}\,\big|\,\widetilde{B}_{r}\big)\geq\sum_{j=h_{r}+1}^{h_{r-1}}c_{2}n\cdot\frac{c_{1}(j-h_{r})}{2n}=\frac{c_{1}c_{2}}{4}\lfloor\theta n\rfloor(\lfloor\theta n\rfloor-1)\geq c_{3}\theta^{2}n^{2}.$ (4.19)
@@
$\begin{aligned} \mathbb{E}_{n}\big((\widetilde{\mathcal{N}}_{i,r})^{2}\,\big|\,\widetilde{B}_{r}\big)&=\mathbb{E}_{n}\big(\widetilde{\mathcal{N}}_{i,r}\,\big|\,\widetilde{B}_{r}\big)+\sum_{(j,l)\neq(j^{\prime},l^{\prime})}\mathbb{P}_{n}\big(\tau_{i}(y^{j}_{l}),\,\tau_{i}(y^{j^{\prime}}_{l^{\prime}})<\infty\,\big|\,\widetilde{B}_{r}\big)\\ &\leq\mathbb{E}_{n}\big(\widetilde{\mathcal{N}}_{i,r}\,\big|\,\widetilde{B}_{r}\big)+2\sum_{(j,l)\neq(j^{\prime},l^{\prime}):j\leq j^{\prime}}\mathbb{P}_{n}\big(\tau_{i}(y^{j}_{l}),\,\tau_{i}(y^{j^{\prime}}_{l^{\prime}})<\infty\,\big|\,\widetilde{B}_{r}\big),\end{aligned}$ (4.20)
@@
By (4.24), (H1) and (H2), for $(j,l)\neq(j^{\prime},l^{\prime})$ with $j\leq j^{\prime}$,
${\mathbb{P}}_{n}\big(\tau_{i}(v^{j}_{l})<\infty\,\big|\,\tau_{i}(v^{j^{\prime}}_{l^{\prime}})<\infty,\,\widetilde{B}_{r}\big)$ $\leq\frac{1}{m}\bigg\{1-\big(1+o_{n}(1)\big)\bigg[1-\frac{c_{4}m}{n}\bigg]^{\theta n}\bigg\}\leq c_{5}\theta,$
@@
$\leq c_{5}\theta\cdot\sum_{(j,l)\neq(j^{\prime},l^{\prime}):j\leq j^{\prime}}\mathbb{P}_{n}\big(\tau_{i}(y^{j^{\prime}}_{l^{\prime}})<\infty\,\big|\,\widetilde{B}_{r}\big)\leq c_{6}\theta^{2}n^{2}\cdot\mathbb{E}_{n}\big(\widetilde{\mathcal{N}}_{i,r}\,\big|\,\widetilde{B}_{r}\big).$
Substituting this estimate into (4.20) yields that for all large $n$,
@@
Finally, taking $\gamma=\tfrac{1}{2}c_{3}\theta^{2}$, we obtain
@@
This completes the proof. ∎
@@
$\mathbb{P}_{n}\big(\tau(v^{j}_{l})=\infty\,\big|\,\widetilde{B}_{r}\big)=\prod_{h=0}^{j-h_{r}-1}\Bigg[1-\frac{m\sigma^{2}_{n}(j-h)}{q_{n}(j-h)}\big(1+o_{n}(1)\big)\Bigg].$ (4.21)
Assuming Lemma 4.15, the desired estimate (4.16) follows directly from the exchangeability of the vector $\big(\mathscr{T}^{n}(v):|v|=h_{r}\big)$. ∎
Proof of Lemma 4.15.
@@
$\mathbb{P}_{n}\big(\tau(v^{j}_{l})>h+1\,\big|\,\tau(v^{j}_{l})>h,\,\widetilde{B}_{r}\big)=1-\frac{m\sigma^{2}_{n}(j-h)}{q_{n}(j-h)}\big(1+o_{n}(1)\big).$ (4.22)
@@
Then we have the decomposition $\widetilde{B}_{r}=\widetilde{B}^{[j-h,h_{r-1}]}_{r}\cap\widetilde{B}^{[h_{r},j-h]}_{r}$. By Proposition 3.8 and the uniform ordering of vertices $v^{j}_{l}$, the vector
@@
together with the event $\{\tau(v^{j}_{l})>h\}\cap\widetilde{B}^{[j-h,h_{r-1}]}_{r}$, forms a (m,jh,m+1)(m,j-h,m+1)-uniform selection pair, with $h_{r-1}$ serving as the base height $h^{*}$ in Definition 4.11. Hence, by Lemma 4.13,
$\begin{aligned} \mathbb{P}_{n}\big(\tau(v^{j}_{l})>h+1,\,\widetilde{B}_{r}\big)&=\mathbb{P}_{n}\big(\tau(v^{j}_{l})>h,\,\widetilde{B}^{[j-h,h_{r-1}]}_{r}\big)\\ &\quad\cdot\mathbb{P}_{n}\Big[\mathfrak{p}(u^{j-h}_{1}),\ldots,\mathfrak{p}(u^{j-h}_{m+1})\text{ are distinct, }\#\big\{\text{Anc}(u^{j-h}_{i},h_{r}):1\leq i\leq m\big\}=m\Big].\end{aligned}$ (4.23)
@@
$\Big(\{\mathfrak{p}(u^{j-h}_{1}),\ldots,\mathfrak{p}(u^{j-h}_{m+1})\text{ are distinct}\},\ \big\{\mathfrak{p}(u^{j-h}_{j}):1\leq i\leq m\big\}\Big)$
is a $(m+1,1,m)$-uniform selection pair with jhj-h as the base height. Applying Lemma 4.13 again yields